I'm completely lost...
How do you factor these?
3x^2+5x+2
5x^2-14x-3
4x^2+11x+6
6x^2+17x+10
Thank you!

Answer by Edwin McCravy(20059)



3x²+5x+2

Multiply the 3 by the 2 ignoring signs.  Get 6.

Write down all the ways to have two positive integers
which have product 6, starting with 6*1

6*1
3*2

Since the last sign in 3x²+5x+2 is +, ADD them,
and place the SUM out beside that:

6*1    6+1=7
3*2    3+2=5

Now, again ignoring signs, we find in that list of
sums the coefficient of the middle term in 3x²+5x+2
which is 5

So we replace the number 5 by 3+2

3x²+5x+2
3x²+(3+2)x+2 

Then we distribute to remove the parentheses:

3x²+3x+2x+2

Factor the first two terms 3x²+3x by taking out the
greatest common factor, getting 3x(x+1)

Factor the last two terms +2x+2 by taking out the
greatest common factor, +2, getting +2(x+1)

So we have

3x(x+1)+2(x+1)

Notice that there is a common factor, (x+1)

3x(x+1)+2(x+1)

which we can factor out leaving the 3x and the +2 to put 
in parentheses:

(x+1)(3x+2)

------------------------------

5x²-14x-3
 
Multiply the 5 by the 3 ignoring signs.  Get 15

Write down all the ways to have two positive integers
which have product 15, starting with 15*1

15*1
 5*3

Since the last sign in 5x²-14x-3 is -, SUBTRACT them,
and place the DIFFERENCE out beside that:

15*1   15-1=14
 5*3    5-3=2

Now, again ignoring signs, we find in that list of
sums the coefficient of the middle term in 5x²-14x-3

So we replace the number 14 by 15-1

5x²-14x-3
5x²-(15-1)x-3

Then we distribute to remove the parentheses:

5x²-15x+1x-3

Factor the first two terms 5x²-15x by taking out the
greatest common factor, 5x, getting 5x(x-3)

Factor the last two terms +1x-3 by taking out the
greatest common factor, +1, getting +1(x-3)

So we have

5x(x-3)+1(x-3)

Notice that there is a common factor, (x-3)

5x(x-3)+1(x-3)

which we can factor out leaving the 5x and the +1 to put 
in parentheses:

(x-3)(5x+1)

---------------------------

The others are done just like the first one.

I wrote a lesson on this method of factoring.  Go here:

http://www.algebra.com/algebra/homework/Expressions-with-variables/change-this-name32371.lesson




Edwin